OpenStudy (anonymous):

Which of the following systems of equations has infinitely many solutions? 3x − 3y = 11; 6y = 9x − 22 5y = -9x − 6; -27x + 15y = 15 4x − 7y = 7; -14y = 8x − 14 -4x + 2 = 3y; 12y − 8 = -16x

OpenStudy (anonymous):

The fourth one because an identity creates infinitely many solutions. -4(-4x - 3y = -2) -16x - 12y = -8 0 = 0 This is an identity.